# Fast Policy Learning for Linear-Quadratic Control with Entropy Regularization

### v3

@@ 5.2 Super-linear Local Convergence
$\kappa:=$ $\frac{\rho+\|A\|}{|\sigma_{\min}(B)|},\quad c:=2\rho\xi_{\gamma\rho}\|B\|(\|Q\|+\|R\|\kappa^{2})+\frac{1}{\mu}\|S^{*}\|\|R\|\cdot(\kappa+\|K^{*}\|),$ (27)
$c_{1}:=$ $\left(\xi_{\gamma,\rho}\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\zeta_{\gamma,\rho}\|B\Sigma^{*}B^{\top}+W\|\right)\cdot 2\rho\|B\|$
@@
$c_{2}:=$ $\frac{c\tau\gamma\omega_{\gamma,\rho}\|B\|^{4}}{2\sigma_{\min}(R)^{2}}.$
Note that $\kappa\geq\|K\|$ for any admissible $K$, since $\|K\|\leq\frac{\|BK\|}{|\sigma_{\min}(B)|}\leq\frac{\|A-BK\|+\|A\|}{|\sigma_{\min}(B)|}\leq\frac{\rho+\|A\|}{|\sigma_{\min}(B)|}=\kappa,$ for any $K\in\Omega.$
@@
Theorem 5.3 (Super-linear local convergence of (IPO)).
Let $c_{1}$ and $c_{2}$ be defined in (27). Let $\delta:=\min\left\{\frac{1}{c_{1}+c_{2}}\sigma_{\min}(S^{*}),\frac{\rho-\|A-BK^{*}\|}{\|B\|}\right\}$. Suppose that the initial policy $(K^{(0)},\Sigma^{(0)})$ satisfies
$C(K^{(0)},\Sigma^{(0)})-C(K^{*},\Sigma^{*})\leq\left(\frac{1}{\mu}-\frac{1}{\|S^{*}\|}\right)^{-1}\sigma_{\min}\left(R+\gamma BP_{K^{*}}B\right)\delta^{2},$ (28)
@@
The following Lemma 5.4 is critical for establishing this super-linear local convergence: it shows that there is a contraction if the differences between two discounted state correlation matrices $\|S^{(t+1)}-S^{*}\|$ is small enough. Then, by the perturbation analysis for $S_{K,\Sigma}$ (Lemma 5.5), one can bound $\|S^{(t+1)}-S^{*}\|$ by $\|K^{(t)}-K^{*}\|$ (Lemma 5.6). The proof of Theorem 5.3 follows by ensuring the admissibility of model parameters $\left\{K^{(t)},\Sigma^{(t)}\right\}_{t=0}^{\infty}$ along all the updates.
@@
Suppose that the updating rule (IPO) satisfies $\|S^{(t+1)}-S^{*}\|\leq\sigma_{\min}(S^{*})$ for all $t\geq 0,$ then
@@
Assume that the update of parameter $(K,\Sigma)$ follows the updating rule in (IPO). Let $c_{1},c_{2}$ be defined in (27). Then for $K^{(t+1)}$ satisfying $\|A-BK^{(t+1)}\|\leq\rho$, we have
$\|S^{(t+1)}-S^{*}\|\leq(c_{1}+c_{2})\|K^{(t)}-K^{*}\|.$
Proof.
(of Theorem 5.3). Fix integer $t\geq 0$. Observe that
$\left(1-\frac{\mu}{\|S_{K^{*},\Sigma^{*}}\|}\right)\left(C(K^{(t)},\Sigma^{(t)})-C(K^{*},\Sigma^{*})\right)\overset{(a)}{\geq}C(K^{(t+1)},\Sigma^{(t+1)})-C(K^{*},\Sigma^{*})$ (29)
$\overset{(b)}{\geq}$ $\mu\sigma_{\min}\left(R+\gamma BP_{K^{*}}B\right)\|K^{(t+1)}-K^{*}\|^{2}+f_{K^{*}}\left(\Sigma^{*}\right)-f_{K^{*}}\left(\Sigma^{(t+1)}\right)$
@@
$(a)$ is clear from the contraction property in Lemma 5.2; $(b)$ follows from Lemma 3.3 and (64); to obtain $(c)$, note that $f_{K^{*}}\left(\Sigma^{*}\right)-f_{K^{*}}\left(\Sigma^{(t+1)}\right)\geq 0,$ since $\Sigma^{*}$ is the maximizer of $f_{K^{*}}.$ Thus, (29) and (28) imply
$\|K^{(t+1)}-K^{*}\|\leq\min\left\{\frac{1}{c_{1}+c_{2}}\sigma_{\min}(S^{*}),\frac{\rho-\|A-BK^{*}\|}{\|B\|}\right\},$
which suggests that
$\|A-BK^{(t+1)}\|\leq\|A-BK^{*}\|+\|B\|\|K^{(t+1)}-K^{*}\|\leq\rho.$
The above inequality along with Lemma 5.6 leads
$\|S^{(t+1)}-S^{*}\|\leq(c_{1}+c_{2})\|K^{(t)}-K^{*}\|\leq\sigma_{\min}(S^{*}).$ (30)
Then, apply the Lemma 5.4 and (30) to get:
$C(K^{(t+1)},\Sigma^{(t+1)})-C(K^{*},\Sigma^{*})$ (31)
$\leq\frac{(c_{1}+c_{2})}{\sigma_{\min}(S^{*})}\|K^{(t)}-K^{*}\|\left(C(K^{(t)},\Sigma^{(t)})-C(K^{*},\Sigma^{*})\right).$
Using the same reasoning as in (29) $(a)$ to $(c)$, we have
$C(K^{(t)},\Sigma^{(t)})-C(K^{*},\Sigma^{*}){\geq}$ $\mu\sigma_{\min}\left(R+\gamma B^{\top}P_{K^{*}}B\right)\|K^{(t)}-K^{*}\|^{2}.$
Plugging the above results in (31) finishes the proof. ∎
6 Application: Transfer Learning
@@
Problem set-up and main results.
We analyze two environments $\mathcal{M}:=(A,B)$ and ¯$\overline{\mathcal{M}}:=(\overline{A},\overline{B})$, with $(K^{*},\Sigma^{*})$ and $(\overline{K}^{*},\overline{\Sigma}^{*})$ as their respective optimal policies. Assume that one has access to the optimal (regularized) policy $(K^{*},\Sigma^{*})$ for environment $\mathcal{M}$, called the well-understood environment. We use $(K^{*},\Sigma^{*})$ as a policy initialization for the less understood environment $\overline{\mathcal{M}}$, called the new environment. The goal is to investigate under what conditions this initialization enters the super-linear convergence regime of $\overline{\mathcal{M}}$.
[…]
$\|A\|:=\|A\|_{q}=\sup_{x\neq 0}\left\{\frac{\|Ax\|_{q}}{\|x\|_{q}},\,\,x\in\mathbb{R}^{n_{2}}\right\}.$ (32)
For ease of the analysis and to make the two environments comparable, the following assumptions are made:
@@
Similar to $P_{K}$ defined in (17) for environment $\mathcal{M}$, let us define the Riccati equation for the new environment $\overline{M}$ as:
$\overline{P}_{K}$ $=$ $\gamma(\overline{A}-\overline{B}K)^{\top}\overline{P}_{K}(\overline{A}-\overline{B}K)+Q+K^{\top}RK,$ (33)
and define $\kappa^{\prime},c^{\prime},c_{1}^{\prime},c_{2}^{\prime}$ in the same way as (27) with $(\overline{A},\overline{B})$ replacing $(A,B)$.
@@
Let $\delta^{\prime}:=\min\left\{\frac{1}{c^{\prime}_{1}+c^{\prime}_{2}}\sigma_{\min}(S_{\overline{K}^{*},\overline{\Sigma}^{*}}),\frac{\rho-\|\overline{A}-\overline{B}\overline{K}^{*}\|}{\|\overline{B}\|}\right\}$ and $c_{\gamma,\rho}:=\max\{\frac{\gamma}{1-\gamma},\frac{\gamma\rho}{1-\gamma\rho^{2}}\}$. If the following condition is satisfied:
$\Big{(}\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big{)}\leq\frac{\left(\frac{1}{\mu}-\frac{1}{\|S^{*}\|}\right)^{-1}\sigma_{\min}\left(R+\gamma\overline{B}\overline{P}_{\overline{K}^{*}}\overline{B}\right)(\delta^{\prime})^{2}}{4c_{\gamma,\rho}\Big{(}\left\|\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}x_{0}^{\top}]+W\right\|_{q}+\frac{\gamma}{1-\gamma}+1\Big{)}\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}},$ (34)
then $(K^{*},\Sigma^{*})$ is within the super-linear convergence region of environment $\overline{\mathcal{M}}$, i.e., for all $t\geq 0$,
$C^{\overline{\mathcal{M}}}(\overline{K}^{(t+1)},\overline{\Sigma}^{(t+1)})-C^{\overline{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})\leq\frac{(c^{\prime}_{1}+c^{\prime}_{2})\cdot\left(C^{\overline{\mathcal{M}}}(\overline{K}^{(t)},\overline{\Sigma}^{(t)})-C^{\overline{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})\right)^{1.5}}{\sigma_{\min}(S_{\overline{K}^{*},\overline{\Sigma}^{*}})\sqrt{\mu\sigma_{\min}(R+\gamma\overline{B}^{\top}\overline{P}_{\overline{K}^{*}}\overline{B})}},$
if the initial policy follows $(\overline{K}^{(0)},\overline{\Sigma}^{(0)})=(K^{*},\Sigma^{*})$ and the policy updates according to (IPO).
Proof.
@@
$\left|C^{\overline{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})-C^{\overline{\mathcal{M}}}(K^{*},\Sigma^{*})\right|\leq$ $\left|C^{\overline{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})-C^{{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})\right|$ (35)
$+\left|C^{\overline{\mathcal{M}}}(K^{*},\Sigma^{*})-C^{{\mathcal{M}}}(K^{*},\Sigma^{*})\right|.$
For any given policy $(K,\Sigma)$ that is admissible to both $\mathcal{M}$ and $\overline{\mathcal{M}}$, we have $J^{\mathcal{M}}_{K,\Sigma}(x)=x^{\top}P_{K}x+q_{K,\Sigma}$ and $J^{\overline{\mathcal{M}}}_{K,\Sigma}(x)=x^{\top}\overline{P}_{{K}}x+\overline{q}_{K,\Sigma}$ with some symmetric positive definite matrices $P_{K},\overline{P}_{K}\in\mathbb{R}^{n\times n}$ satisfying (17) and (33) respectively. In addition, the constants $q_{K,\Sigma}$ takes the form of (17) and $\overline{q}_{K,\Sigma}\in\mathbb{R}$ takes the following form
$\overline{q}_{K,\Sigma}=$ $\frac{1}{1-\gamma}\Big{[}-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma\right)\right)+\operatorname{Tr}\left(\Sigma R+\gamma(\Sigma(\overline{B})^{\top}\overline{P}_{K}\overline{B}+W\overline{P}_{K})\right)\Big{]}.$
We can see that
@@
$+\gamma(\overline{A}-\overline{B}K)^{\top}(P_{K}-\overline{P}_{K})(\overline{A}-\overline{B}K).$
@@
$\|P_{K}-\overline{P}_{K}\|\leq\frac{2\gamma\rho}{1-\gamma\rho^{2}}(\|A-\overline{A}\|+\|K\|\,\|B-\overline{B}\|)\|P_{K}\|.$ (36)
@@
$\overline{q}_{K,\Sigma}-q_{K,\Sigma}$ $=$ $\frac{\gamma}{1-\gamma}\Biggr{(}\text{Tr}\Big{(}(B-\overline{B})^{\top}P_{K}\,B\Big{)}+\text{Tr}\Big{(}\overline{B}^{\top}P_{K}\,(B-\overline{B})\Big{)}$
$\quad\quad\quad\quad+\text{Tr}\Big{(}\overline{B}^{\top}(P_{K}-\overline{P}_{K})\,\overline{B}\Big{)}+\text{Tr}(W(P_{K}-\overline{P}_{K}))\Biggr{)}.$
@@
$\left|C^{\overline{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})-C^{{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})\right|$ (37)
$\leq$ $\Big{|}\text{Tr}\Big{(}(P_{\overline{K}^{*}}-\overline{P}_{\overline{K}^{*}})(D_{0}+W)\Big{)}\Big{|}+\frac{\gamma}{1-\gamma}\left|\text{Tr}\Big{(}(B-\overline{B})^{\top}P_{K^{*}}\,B\Big{)}\right|$
@@
where the last inequality follows from $|\text{Tr}(AB)|\leq\|A^{\prime}\|_{p}\,\|B\|_{q}$ when $11/p+1/q=$ 1. This is a consequence of combining von Neumann’s trace inequality with Hölder’s inequality for Euclidean space.
(17) suggests that, for any admissible $K$ such that $\|K\|\leq 1$, we have $\|P_{K}\|\leq\gamma\rho^{2}\|P_{K}\|+\|Q\|+\|R\|.$ Hence, $\|P_{K}\|\leq\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}$. Combining this bound with (36) and the fact that $\|B\|_{q},\|\overline{B}\|_{q},\|K^{*}\|\leq 1$, (37) can be bounded such that
[…]
$+\frac{2\gamma}{1-\gamma}\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\|B-\overline{B}\|_{q}.$
[…]
(38)$\leq 2c_{\gamma,\rho}\Big{(}\|D_{0}+W\|_{q}+\frac{\gamma}{1-\gamma}+1\Big{)}\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\Big{(}\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big{)}.$ ) . (39)

### v4

@@ 5.2 Local Super-linear Convergence
(5.2) $\kappa$= $\frac{\rho+\|A\|}{|\sigma_{\min}(B)|},\quad c=2\rho\xi_{\gamma\rho}\|B\|(\|Q\|+\|R\|\kappa^{2})+\frac{1}{\mu}\|S^{*}\|\|R\|\cdot(\kappa+\|K^{*}\|),$
$c_{1}$= $\left(\xi_{\gamma,\rho}\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\zeta_{\gamma,\rho}\|B\Sigma^{*}B^{\top}+W\|\right)\cdot 2\rho\|B\|$
@@
$c_{2}$= $\frac{c\tau\gamma\omega_{\gamma,\rho}\|B\|^{4}}{2\sigma_{\min}(R)^{2}}.$
Note that for any $K\in\Omega$, $\|K\|\leq\frac{\|BK\|}{|\sigma_{\min}(B)|}\leq\frac{\|A-BK\|+\|A\|}{|\sigma_{\min}(B)|}\leq\frac{\rho+\|A\|}{|\sigma_{\min}(B)|}=\kappa.$
@@
Theorem 5.3 (Local super-linear convergence of (IPO)).
Let $c_{1}$ and $c_{2}$ be defined in (5.2). Let $\delta:=\min\left\{\frac{1}{c_{1}+c_{2}}\sigma_{\min}(S^{*}),\frac{\rho-\|A-BK^{*}\|}{\|B\|}\right\}$. Suppose that the initial policy $(K^{(0)},\Sigma^{(0)})$ satisfies
(5.3) $C(K^{(0)},\Sigma^{(0)})-C(K^{*},\Sigma^{*})\leq\left(\frac{1}{\mu}-\frac{1}{\|S^{*}\|}\right)^{-1}\sigma_{\min}\left(R+\gamma BP_{K^{*}}B\right)\delta^{2},$
@@
The following Lemma 5.4 is critical for establishing this local super-linear convergence: it shows that there is a contraction if the differences between two discounted state correlation matrices $\|S^{(t+1)}-S^{*}\|$ is small enough. Then, by the perturbation analysis for $S_{K,\Sigma}$ (Lemma 5.5), one can bound $\|S^{(t+1)}-S^{*}\|$ by $\|K^{(t)}-K^{*}\|$ (Lemma 5.6). The proof of Theorem 5.3 follows by ensuring the admissibility of model parameters $\left\{K^{(t)},\Sigma^{(t)}\right\}_{t=0}^{\infty}$ along all the updates.
@@
Suppose that $\|S^{(t+1)}-S^{*}\|\leq\sigma_{\min}(S^{*})$ for all $t\geq 0$ when updating with (IPO), then
@@
Assume that the update of parameter $(K,\Sigma)$ follows (IPO). Let $c_{1},c_{2}$ be defined in (5.2). Then for $K^{(t+1)}$ satisfying $\|A-BK^{(t+1)}\|\leq\rho$, we have $\|S^{(t+1)}-S^{*}\|\leq(c_{1}+c_{2})\|K^{(t)}-K^{*}\|.$
Proof 5.7.
(of Theorem 5.3). First, Theorem 2.1 shows that for an optimal $K^{*}$, $K^{*}=\gamma(R+\gamma B^{\top}P_{K^{*}}B)^{-1}B^{\top}P_{K^{*}}A.$ Then, by the definition of $E_{K}$ in Lemma 3.1,
(5.4) $E_{K^{*}}=-\gamma B^{\top}P_{K^{*}}A+(\gamma B^{\top}P_{K^{*}}B+R)K^{*}=0.$
Fix integer $t\geq 0$. Observe that
(5.5) $\left(1-\frac{\mu}{\|S_{K^{*},\Sigma^{*}}\|}\right)\left(C(K^{(t)},\Sigma^{(t)})-C(K^{*},\Sigma^{*})\right)\overset{(a)}{\geq}C(K^{(t+1)},\Sigma^{(t+1)})-C(K^{*},\Sigma^{*})$
$\overset{(b)}{\geq}\mu\sigma_{\min}\left(R+\gamma BP_{K^{*}}B\right)\|K^{(t+1)}-K^{*}\|^{2}+f_{K^{*}}(\Sigma^{*})-f_{K^{*}}(\Sigma^{(t+1)})$
@@
$(a)$ is from the contraction property in Lemma 5.2; $(b)$ follows from Lemma 3.3 and (5.4); to obtain $(c)$, note that $f_{K^{*}}\left(\Sigma^{*}\right)-f_{K^{*}}\left(\Sigma^{(t+1)}\right)\geq 0,$ since $\Sigma^{*}$ is the maximizer of $f_{K^{*}}.$ Thus, (5.5) and (5.3) imply $\|K^{(t+1)}-K^{*}\|\leq\delta$ which suggests that $\|A-BK^{(t+1)}\|\leq\|A-BK^{*}\|+\|B\|\|K^{(t+1)}-K^{*}\|\leq\rho.$ Then by Lemma 5.6,
(5.6) $\|S^{(t+1)}-S^{*}\|\leq(c_{1}+c_{2})\|K^{(t)}-K^{*}\|\leq\sigma_{\min}(S^{*}).$
Thus, one can apply Lemma 5.4 to get:
(5.7) $C(K^{(t+1)},\Sigma^{(t+1)})-C(K^{*},\Sigma^{*})\leq\frac{c_{1}+c_{2}}{\sigma_{\min}(S^{*})}\|K^{(t)}-K^{*}\|(C(K^{(t)},\Sigma^{(t)})-C(K^{*},\Sigma^{*})).$
Using the same reasoning as in (5.5) $(a)$ to $(c)$, we have $C(K^{(t)},\Sigma^{(t)})-C(K^{*},\Sigma^{*}){\geq}\\ \mu\sigma_{\min}\left(R+\gamma B^{\top}P_{K^{*}}B\right)\|K^{(t)}-K^{*}\|^{2}$ and plug it in (5.7) finishes the proof.
6 Transfer Learning for RL
@@
Problem set-up and main results
We analyze two environments $\mathcal{M}:=(A,B)$ and $\overline{\mathcal{M}}:=(\overline{A},\overline{B})$, with $(K^{*},\Sigma^{*})$ and $(\overline{K}^{*},\overline{\Sigma}^{*})$ as their respective optimal policies and $C$ and $\overline{C}$ as their respective objective functions. Assume that one has access to the optimal (regularized) policy $(K^{*},\Sigma^{*})$ for environment $\mathcal{M}$, called the well-understood environment. We use $(K^{*},\Sigma^{*})$ as a policy initialization for the less understood environment $\overline{\mathcal{M}}$, called the new environment. The goal is to investigate under what conditions this initialization enters the super-linear convergence regime of $\overline{\mathcal{M}}$.
Throughout this section, we specify the operator norm $\|\cdot\|$ as the one associated with vector $q$-norm. Namely, for $q\in(0,1)$ and $A\in\mathbb{R}^{n_{1}\times n_{2}}$ for some positive integers $n_{1},n_{2}$: $\|A\|:=\|A\|_{q}=\sup_{x\neq 0}\left\{\frac{\|Ax\|_{q}}{\|x\|_{q}},\,\,x\in\mathbb{R}^{n_{2}}\right\}.$ For ease of the analysis and to make the two environments comparable, the following assumptions are made:
@@
Similar to $P_{K}$ defined in (3.5) for environment $\mathcal{M}$, let us define the Riccati equation for the new environment $\overline{M}$ as:
(6.1) $\overline{P}_{K}$ $=$ $\gamma(\overline{A}-\overline{B}K)^{\top}\overline{P}_{K}(\overline{A}-\overline{B}K)+Q+K^{\top}RK,$
and define $\kappa^{\prime},c^{\prime},c_{1}^{\prime},c_{2}^{\prime}$ in the same way as (5.2) with $(\overline{A},\overline{B})$ replacing $(A,B)$.
The following theorem suggests that if the environments $\mathcal{M}$ and $\overline{\mathcal{M}}$ are sufficiently close in the sense of (6.2), then $(K^{*},\Sigma^{*})$ serves an efficient initial policy for $\overline{\mathcal{M}}$ which directly leads to a super-linear convergence for the new learning problem
@@
Let $c_{\gamma,\rho}:=\max\{\frac{\gamma}{1-\gamma},\frac{\gamma\rho}{1-\gamma\rho^{2}}\}$ and $\delta^{\prime}:=\min\Big\{\frac{1}{c^{\prime}_{1}+c^{\prime}_{2}}\sigma_{\min}(S_{\overline{K}^{*},\overline{\Sigma}^{*}}),\\ \frac{\rho-\|\overline{A}-\overline{B}\overline{K}^{*}\|}{\|\overline{B}\|}\Big\}$. If the following condition is satisfied:
(6.2) $\qquad\Big(\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big)\leq\frac{\left(\frac{1}{\mu}-\frac{1}{\|S^{*}\|}\right)^{-1}\sigma_{\min}\left(R+\gamma\overline{B}\overline{P}_{\overline{K}^{*}}\overline{B}\right)(\delta^{\prime})^{2}}{4c_{\gamma,\rho}\Big(\left\|\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}x_{0}^{\top}]+W\right\|_{q}+\frac{\gamma}{1-\gamma}+1\Big)\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}},$
then $(K^{*},\Sigma^{*})$ is within the super-linear convergence region of environment $\overline{\mathcal{M}}$, i.e.,
$\overline{C}(\overline{K}^{(t+1)},\overline{\Sigma}^{(t+1)})-\overline{C}(\overline{K}^{*},\overline{\Sigma}^{*})\leq\frac{(c^{\prime}_{1}+c^{\prime}_{2})\cdot\left(\overline{C}(\overline{K}^{(t)},\overline{\Sigma}^{(t)})-\overline{C}(\overline{K}^{*},\overline{\Sigma}^{*})\right)^{1.5}}{\sigma_{\min}(S_{\overline{K}^{*},\overline{\Sigma}^{*}})\sqrt{\mu\sigma_{\min}(R+\gamma\overline{B}^{\top}\overline{P}_{\overline{K}^{*}}\overline{B})}},$
for all $t\geq 0$, if the initial policy follows $(\overline{K}^{(0)},\overline{\Sigma}^{(0)})=(K^{*},\Sigma^{*})$ and the policy updates according to (IPO).
Proof 6.2.
@@
(6.3) $\left|\overline{C}(\overline{K}^{*},\overline{\Sigma}^{*})-\overline{C}(K^{*},\Sigma^{*})\right|\leq$ $\left|\overline{C}(\overline{K}^{*},\overline{\Sigma}^{*})-C(\overline{K}^{*},\overline{\Sigma}^{*})\right|+\left|\overline{C}(K^{*},\Sigma^{*})-C(K^{*},\Sigma^{*})\right|.$
For any given policy $(K,\Sigma)$ that is admissible to both $\mathcal{M}$ and $\overline{\mathcal{M}}$, we have $J_{K,\Sigma}(x)=x^{\top}P_{K}x+q_{K,\Sigma}$ and $\overline{J}_{K,\Sigma}(x)=x^{\top}\overline{P}_{{K}}x+\overline{q}_{K,\Sigma}$ with some symmetric positive definite matrices $P_{K},\overline{P}_{K}\in\mathbb{R}^{n\times n}$ satisfying (3.5) and (6.1) respectively. In addition, the constants $q_{K,\Sigma}$ takes the form of (3.5) and $\overline{q}_{K,\Sigma}\in\mathbb{R}$ takes the form of $\overline{q}_{K,\Sigma}=\frac{1}{1-\gamma}\Big(-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma\right)\right)+\operatorname{Tr}\left(\Sigma R+\gamma(\Sigma(\overline{B})^{\top}\overline{P}_{K}\overline{B}+W\overline{P}_{K})\right)\Big).$ Note that
@@
$\quad+\gamma(\overline{A}-\overline{B}K)^{\top}(P_{K}-\overline{P}_{K})(\overline{A}-\overline{B}K).$
@@
(6.4) $\|P_{K}-\overline{P}_{K}\|\leq\frac{2\gamma\rho}{1-\gamma\rho^{2}}(\|A-\overline{A}\|+\|K\|\,\|B-\overline{B}\|)\|P_{K}\|.$
@@
$\overline{q}_{K,\Sigma}-q_{K,\Sigma}$ $=$ $\frac{\gamma}{1-\gamma}\Big(\text{Tr}\Big((B-\overline{B})^{\top}P_{K}\,B\Big)+\text{Tr}\Big(\overline{B}^{\top}P_{K}\,(B-\overline{B})\Big)$
$\quad\quad\quad\quad+\text{Tr}\Big(\overline{B}^{\top}(P_{K}-\overline{P}_{K})\,\overline{B}\Big)+\text{Tr}(W(P_{K}-\overline{P}_{K}))\Big).$
@@
(6.5) $\left|\overline{C}(\overline{K}^{*},\overline{\Sigma}^{*})-C(\overline{K}^{*},\overline{\Sigma}^{*})\right|$
$\leq$ $\Big|\operatorname{Tr}\Big((P_{\overline{K}^{*}}-\overline{P}_{\overline{K}^{*}})(D_{0}+W)\Big)\Big|+\frac{\gamma}{1-\gamma}\left|\text{Tr}\Big((B-\overline{B})^{\top}P_{K^{*}}\,B\Big)\right|$
@@
where the last inequality follows from $|\text{Tr}(AB)|\leq\|A^{\prime}\|_{p}\,\|B\|_{q}$ when $1/p+1/q=1$. This is a consequence of combining von Neumann’s trace inequality with Hölder’s inequality for Euclidean space.
(3.5) suggests that, for any admissible $K$ such that $\|K\|\leq 1$, we have $\|P_{K}\|\leq\gamma\rho^{2}\|P_{K}\|+\|Q\|+\|R\|.$ Hence, $\|P_{K}\|\leq\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}$. Combining this bound with (6.4) and the fact that $\|B\|_{q},\|\overline{B}\|_{q},\|K^{*}\|\leq 1$, (6.5) can be bounded such that
(6.5) $\leq\Big(\|D_{0}+W\|_{q}+\frac{\gamma}{1-\gamma}\Big)\frac{2\gamma\rho}{1-\gamma\rho^{2}}\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\Big(\|A-\overline{A}\|+\|B-\overline{B}\|\Big)$
$+\frac{2\gamma}{1-\gamma}\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\|B-\overline{B}\|_{q}$
(6.6) $\leq 2c_{\gamma,\rho}\Big(\|D_{0}+W\|_{q}+\frac{\gamma}{1-\gamma}+1\Big)\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\Big(\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big).$